OpenStudy (anonymous):

What is the shortest possible distance from a point on the parabola y=x^2 to the point (0,1)?

OpenStudy (amistre64):

a perpendicular line

OpenStudy (amistre64):

and wont it be 2 points?

OpenStudy (amistre64):

distance, ok, gotta learn to read

OpenStudy (anonymous):

still can use that though right?

OpenStudy (amistre64):

(x,x^2) -(0, 1 ) ---------- x,x^2-1 distance = sqrt((x^2)+(x^2-1)^2)

OpenStudy (amistre64):

id prooly try to get the point thru calculus and then find the distance

OpenStudy (anonymous):

make your life easy and ignore the radical.

OpenStudy (amistre64):

dD/dx = 4x^3 - 2x = 2x (2x^2 - 1) = 0 x = 0, x^2 = 1/2 then right?

OpenStudy (anonymous):

yeah i am getting a different answer

OpenStudy (anonymous):

i get \[x=\pm\frac{\sqrt{2}}{2}\]

OpenStudy (amistre64):

same here; or rather the shortest distance is about; .866...

OpenStudy (jamesj):

and x = 0 is a local max

OpenStudy (jamesj):

sorry 'bout that.

OpenStudy (anonymous):

here is an alternate method, not that the other one doesn't work. you get the slope as \[\frac{x^2-1}{x}\] and it must be perpendicular to the tangent line at \[y=x^2\] so you must have \[\frac{x^2-1}{x}=-\frac{1}{2x}\] solve for x and get \[x=\pm\frac{\sqrt{2}}{2}\]

myininaya (myininaya):

\[d=\sqrt{(x-0)^2+(x^2-1)^2}\] \[d^2=x^2+(x^2-1)^2 \] \[(d^2)'=2x+2(x^2-1)(2x) =(2x)(1+2(x^2-1))=2x(2x^2-2+1)=2x(2x^2-1)\] \[=> x=0, x=\frac{1}{\sqrt{2}}, x=\frac{-1}{\sqrt{2}}\]

myininaya (myininaya):

and we need to check to see they are max or mins

OpenStudy (amistre64):

at x=1 the distance is obviously 1 :)

OpenStudy (amistre64):

err... at x=0

OpenStudy (anonymous):

using my method (he boasts) i don't need to check

OpenStudy (jamesj):

Well, D'' = 12x^2 - 2 hence D''(0) = -2 < 0 and hence x = 0 is a local max.

OpenStudy (amistre64):

\[f(x)=x^2;\ f'(x)=2x;\ perp.slope=-\frac{1}{2x}\] \[y = -\frac{1(x)}{2x}+\frac{1(0)}{2x}+1\] \[y = -\frac{1}{2}+1\] \[y = \frac{1}{2};\ x^2=\frac{1}{2};\ x=\frac{1}{\sqrt{2}}\]